OpenStudy (anonymous):

Write the equation in slope-intercept form: 5x - 3y = 12.

OpenStudy (anonymous):

Just do your algebra to isolate y.

OpenStudy (anonymous):

-5x from both sides give -3y = -5x + 12 / -3 on both sides gives y = 5/3 * x + 4

OpenStudy (angela210793):

-3y=12-5x 3y=5x-12 y=5x/3-4